# Gravity from entropy

### v2

@@ Abstract
The relation between general relativity, statistical mechanics and information theory is a central research topic in theoretical physics. The interest in the subject has its roots in the discovery that black holes have an entropy [1, 2] and emit Hawking radiation [3]. Recently, important results have been obtained relating information theory, entanglement entropy [4] and gravity [5, 6, 7, 8, 9] involving the holographic principle [10, 11] and beyond [12]. These results define a very active research direction [13] indicating that the quest for an ultimate gravitational theory based on information and statistical mechanics is ongoing. A comprehensive statistical mechanics approach to gravity is expected to give rise to modified Einstein equations that one side can be testable experimentally [14, 15] while on the other side they can bring important conceptual insights into quantum gravity [16, 17, 18, 19, 20, 21].
@@
where here and in the following $\bar{}$ indicates complex conjugation. Let us define the scalar product $|{\bm{\nabla}}\phi|^{2}$ as
@@
Let us now make some remarks about this warm-up derivation. We proposed a statistical mechanics framework that is very inspiring as we get the massless Klein-Gordon equation as the outcome of the minimization of a quantum entropy action for low coupling, i.e. $0<\alpha\ll 1$. However this approach has two important limitations. The first limitation is that the Klein-Gordon equation does not contain the mass term. The second limitation is that in absence of matter fields the metric is not determined. In the following we provide a more comprehensive framework to address these two limitations. This framework can be related to gravity as it gives rise to modified Einstein equations that reduce to the Einstein equations for low coupling.
@@
$\tilde{\bf G}$ $=$ $\tilde{g}+\alpha\Big{(}{D}\ket{\Phi}\bra{\Phi}D+(m+R^{2})\ket{\Phi}\bra{\Phi}\Big{)}$ (32)
@@
where $\alpha,\beta$ are positive constants. According to this choice of the induced metric we obtain that $\tilde{\bf G}$ has the structure given by Eq.(29) with ${G}_{(0)}$ and ${\bf G}_{(1)}$ given by
@@
This action defines a modified theory of gravity. For $\alpha\ll 1$ and $\beta\ll 1$ this action reduces to the Einstein-Hilbert action coupled with the scalar topological field. Indeed in this limit $\bm{\mathcal{L}}$ reduces to
@@
$\delta\bm{\mathcal{L}}$ $=$ $-\mbox{Tr}\tilde{g}{\tilde{\bf G}}^{-1}[\delta{\tilde{\bf G}}\tilde{g}^{-1}+{\tilde{\bf G}}\delta\tilde{g}^{-1}]$ (42)
$=$ $-\mbox{Tr}{\tilde{\bf G}}^{-1}\delta{\tilde{\bf G}}-\mbox{Tr}\tilde{g}\delta\tilde{g}^{-1}$
$=$ $-\mbox{Tr}\delta{\tilde{\bf G}}{\tilde{\bf G}}^{-1}+\mbox{Tr}\delta\tilde{g}\tilde{g}^{-1}.$
@@
In the interesting case in which $\omega_{\mu}=0$, i.e. the bosonic matter field reduced to the scalar $\phi$ the equations of motions reduce to the Klein-Gordon equations with a non-zero mass term, i.e.
@@
As expected the Eq.(45) reduce to the Einstein equations for $\alpha\ll 1$ and $\beta\ll 1$. This implies, among the other things, that by adding to $\tilde{\bf G}$ the term proportional to $\bm{\mathcal{R}}$ we solve the second limitation of the warm-up scenario, and we obtain that the entropic action allows for the full determination of the metric in the vacuum.
@@
In conclusion this work proposes a modified theory of gravity emerging from statistical mechanics and information theory action. The approach is based on the interpretation of the metric as a density matrix albeit its trace is not unitary. The statistical mechanics action is given by the quantum relative entropy between the metric and the metric induced by the matter fields. Here a Lorentzian theory consistent with this fundamental physical interpretation of gravity if formulated. In order to do this we have built the necessary mathematical background to define the entropy and the cross-entropy of rank 2 tensors in a Lorentzian well defined way. A modified gravity is emerging from this framework when the matter fields are described by topological Dirac-Kähler bosons formed by the direct sum between a zero and a one form and the induced metric also depends on the curvature of spacetime. The modified Einstein equations reduce to the Einstein equations in the regime of low coupling.

### v3

@@ Abstract
The relation between general relativity, statistical mechanics and information theory is a central research topic in theoretical physics. The interest in the subject has its roots in the discovery that black holes have an entropy [1, 2] and emit Hawking radiation [3]. Recently, important results have been obtained relating information theory, entanglement entropy [4] and gravity [5, 6, 7, 8, 9] involving the holographic principle [10, 11] and entanglement properties of quantum field theory [12]. These results define a very active research direction [13] indicating that the quest for an ultimate gravitational theory based on information and statistical mechanics is ongoing. A comprehensive statistical mechanics approach to gravity is expected to give rise to modified Einstein equations that one side can be testable experimentally [14, 15] while on the other side they can bring important conceptual insights into quantum gravity [16, 17, 18, 19, 20, 21].
@@
where here and in the following $\bar{}$ indicates complex conjugation. Since ${\bf G}$ should be adimensional it is convenient to work in the units $\hbar=c=1$ and to put $\alpha=\alpha^{\prime}\ell_{P}^{d}$ where $\ell_{P}$ is the Planck length and $\alpha^{\prime}$ is adimensional. Let us define the scalar product $|{\bm{\nabla}}\phi|^{2}$ as
@@
Let us now make some remarks about this warm-up derivation. We proposed a statistical mechanics framework that is very inspiring as we get the massless Klein-Gordon equation as the outcome of the minimization of a quantum entropy action for low coupling, i.e. $0<\alpha^{\prime}\ll 1$. However this approach has two important limitations. The first limitation is that the Klein-Gordon equation does not contain the mass term. The second limitation is that in absence of matter fields the metric is not determined. In the following we provide a more comprehensive framework to address these two limitations. This framework can be related to gravity as it gives rise to modified Einstein equations that reduce to the Einstein equations for low coupling.
@@
$\tilde{\bf G}$ $=$ $\tilde{g}+\alpha\Big{(}{D}\ket{\Phi}\bra{\Phi}D+(m^{2}+R)\ket{\Phi}\bra{\Phi}\Big{)}$ (32)
@@
where $\alpha,\beta$ are positive constants. In particular, since we require $\tilde{\bf G}$ to be adimensional, we need to consider $\alpha=\alpha^{\prime}\ell_{P}^{d}$ and $\beta=\beta^{\prime}\ell_{P}^{2}$ where $\ell_{P}$ is the Planck length and $\alpha^{\prime},\beta^{\prime}$ are adimensional in the units $\hbar=c=1$. According to this choice of the induced metric we obtain that $\tilde{\bf G}$ has the structure given by Eq.(29) with ${G}_{(0)}$ and ${\bf G}_{(1)}$ given by
@@
This action defines a modified theory of gravity. In the linearised limit $\alpha^{\prime}\ll 1$, $\beta^{\prime}\ll 1$ this action reduces to the Einstein-Hilbert action coupled with the scalar topological field. Indeed in this limit $\bm{\mathcal{L}}$ reduces to
@@
$\delta\bm{\mathcal{L}}$ $=$ $-\mbox{Tr}_{M}\tilde{g}{\tilde{\bf G}}^{-1}[\delta{\tilde{\bf G}}\tilde{g}^{-1}+{\tilde{\bf G}}\delta\tilde{g}^{-1}]$ (42)
$=$ $-\mbox{Tr}_{M}{\tilde{\bf G}}^{-1}\delta{\tilde{\bf G}}-\mbox{Tr}_{M}\tilde{g}\delta\tilde{g}^{-1}$
$=$ $-\mbox{Tr}_{M}\delta{\tilde{\bf G}}{\tilde{\bf G}}^{-1}+\mbox{Tr}_{M}\delta\tilde{g}\tilde{g}^{-1}.$
@@
In the interesting case in which $\omega_{\mu}=0$, i.e. the bosonic matter field reduced to the scalar $\phi$ these equations of motions reduce to the Klein-Gordon equations with a non-zero mass term, i.e.
@@
As expected the Eq.(45) reduce to the Einstein equations for $\alpha^{\prime}\ll 1$ and $\beta^{\prime}\ll 1$. This implies, among the other things, that by adding to $\tilde{\bf G}$ the term proportional to $\bm{\mathcal{R}}$ we solve the second limitation of the warm-up scenario, and we obtain that the entropic action allows for the full determination of the metric in the vacuum.
@@
In conclusion this work proposes a modified theory of gravity emerging from statistical mechanics and information theory action. The approach is based on the interpretation of the metric as a density matrix albeit its trace is not unitary, pointing out to possible connections of this approach to von Newmann algebras and the theory of entanglement in quantum field theories [12]. The statistical mechanics action is given by the quantum relative entropy between the metric and the metric induced by the matter fields. Here a Lorentzian theory consistent with this fundamental physical interpretation of gravity is formulated. In order to do this we have built the necessary mathematical background to define the entropy and the cross-entropy of rank 2 tensors in a Lorentzian well defined way. A modified gravity is emerging from this framework when the matter fields are described by topological Dirac-Kähler bosons formed by the direct sum between a zero and a one form and the induced metric also depends on the curvature of spacetime. The modified Einstein equations reduce to the Einstein equations in the regime of low coupling.